86).Selling shirts. If a vendor charges p dollars each for rugby shirts, then he expects to sell 2000 - 100p shirts at a tournament.
a) Find a polynomial R(p) that represents the total revenue when the shirts are p dollars each.
b) Find R(5), R(10), and R(20).
c) Use the bar graph to determine the price that will give the maximum total revenue.

Answer by feliz1965(151)

You were given the function to be 2000 - 100p, where p represents the value in terms of money for each shirt.
Then the polynomial would be
R(p) = 2000 - 100p
To find R(5), R(10) and R(20), replace x with each value individually and simplify the function.
I will find R(5) and you can do the rest.
Let p = 5.
R(5) = 2000 - 100(5)
R(5) = 2000 - 500
R(5) = 1500